OpenStudy (anonymous):

The Express won 80% of the 30 games that they played this season. How many games did they win?

OpenStudy (anonymous):

do u know how to start?

OpenStudy (anonymous):

turn80% to a decimal

OpenStudy (anonymous):

umm..no.. ok see any PERCENT means PER CENT =PER 100

OpenStudy (anonymous):

so 80 percent is 80 per 100 = 80/100 so 80% of 30 = (80/100) X 30 understand?

OpenStudy (anonymous):

\[\huge\frac{ 80 }{ 100}\times 30\]

OpenStudy (anonymous):

100*30/80

OpenStudy (anonymous):

Yeah ... so divide then multiply?

OpenStudy (anonymous):

yes...so can u calculate the answer and check it with me?

OpenStudy (anonymous):

sure thing...hold on for a sec

OpenStudy (anonymous):

Ok so my answer would be 24

OpenStudy (anonymous):

yeah its 24 80*30/100

OpenStudy (anonymous):

aight thanks for your help!!

OpenStudy (anonymous):

yeah thats right!!

OpenStudy (anonymous):

Thanks!!